OpenStudy (anonymous):

Simplify Completely. (9x^-4y^2)^-1/(y^-3z^4) x (3y^-1)^2/(xz^5)

OpenStudy (anonymous):

what is (x^2)^-1?

OpenStudy (anonymous):

(1/x) = (x)^-1

OpenStudy (anonymous):

Sorry this is a question dividing exponents and I thought I could use the ^ symbol to represent that, is that wrong? (9x^-4y^2)^-1/(y^-3z^4) x (3y^-1)^2/(xz^5) means 9x to the -4 power times y to the 2nd power divided by y to the -3 power times z to the 4th power, all in parentheses to the -1st power. then times that by 3y to the -1st power divided by xz to the 5th power in parentheses, all to the 2nd power

OpenStudy (anonymous):

that's not wrong, i meant 1/x equal to x^-1 so, (9x^-4y^2)^-1 = 1/(9x^-4y^2)